OpenStudy (anonymous):

simplify 2(5a-12)=-6(2a-3)+2

OpenStudy (anonymous):

2a-26

OpenStudy (anonymous):

10a-24=-12a+18+2 10a-24=-12a+20 10a+12a=20+24 22a=44 a=44/22 a=2

OpenStudy (anonymous):

-22=a